Suppose a collection of five batteries is connecte

What is the current through the resistor if its value is 2.8 ohms?
Answer #1

a) The emf with greater potential to the left is :

   V = V1+V2-V3+V4-V5

   V = 4.5V + 4.5V - 2.6V + 6.5V - 6.3V

V = 6.6 V

b) Ohm's law says

V = I*R

I = V/R

I = 6.6 V / 2.8 ohms

I = 2.36 A